OpenStudy (anonymous):

1.434 grams of Organic Compound containing Carbon, Hydrogen and Oxygen on Combustion gives 4.4 grams of Carbon dioxide and 2 grams of Water. Find out its Empirical Formula

OpenStudy (australopithecus):

Part 1 Steps related to this type of question: So, since we are getting back a greater mass than what we original burned we can assume this system runs with excess oxygen. So, the first thing we need to do is determine the mass of Carbon in 4.4g of CO2 and the mass of hydrogen in 2g of H2O. To do this apply gravimetric factor. For example if you wanted to figure out the grams of Hydrogen in H2O you would set up the ratio. \[Grams_{H_2O} * \frac{Molecular Mass_{H_2}}{Molecular Mass_{H_2O}} = Grams_{H_2}\] Notice how the units of molecular mass cancel out, as well as the subscript H2O (look at the units as numbers that turn to 1 when divided by each other). Then subtract both of those masses from the 1.434g (the original mass of compound prior to combustion) to determine the mass of oxygen. Part 2 General way to find empirical formula: Once we know the mass of Oxygen, Carbon and Hydrogen convert each mass to moles using: \[Moles = \frac{Grams}{Molecular Mass}\] Next divide the moles determined for oxygen, carbon and hydrogen by the moles of the element with the smallest amount of moles. For example if you had 1mole of oxygen and 2 moles of carbon you would do the following: 1mol/1mol = 1 for oxygen and 2mole/1mole = 2 for carbon these two ratios will give you an empirical formula of CO2. You can see an example of this done here to find: https://openstudy.com/users/australopithecus#/updates/503b5448e4b02c1631bd9522 If you have any questions please feel free to ask

OpenStudy (australopithecus):

ouch I made a mistake it should be 1mol/1mol = 1 for carbon and 2mole/1mole = 2 for oxygen NOT: 1mol/1mol = 1 for oxygen and 2mole/1mole = 2 for carbon if it was this way we would have the formula C2O. One last note if you get numbers with a decimal from the last step multiply (unless the decimal is like 0.99 or 0.98 (in that case just round up) multiply all atoms by an integer to remove the decimal. For example if you have 2moles of carbon and 1.5mole of oxygen you can't have 1.333 atoms of carbon so, multiply it by 3 to make it a whole number 2/1.5 = 1.3333 * 3 = 3.99 = 4 carbon 1.5/1.5 = 1 * 3 = 3 oxygen C4O3